An airplane ticket has a perimeter of 32 centimeters and an area of 60 square centimeters. What are the dimensions of the ticket?
Answer by Murthy(22)

An air plane ticket is rectangular in shape.
Let, its length be x and its breadth be y
its perimeter, p=2(x+y)=32........(1)
Area,A= xy=60............(2)
From,(1) x+y=16....(3)
from, (2) y=60/x...(4)
From (3) & (4) we can form a quadratic equation
x^2-16x+60=0
x=6 cm and 10 cm
y=10 cm and 6 cm
therefore, the dimensions of the ticket are 10 cm and 6 cm.